I need to know how to work out this question and if possible questions like this.


1.sodium chloride is a 2-ion electrolyte dissociating 90% in certain concentrations.... calculate the freezing point of a molal solution

Grade:

1 Answers

Ritwik Garg
12 years ago

For this we use a simple formula given as ΔT=i*kf*m

kf is a constant given as 1.86 for freezing point depression.

i=1+(n-1)*α   here, n=no of ions, α=dissociation degree

ΔT=1.9*1.86*1=3.534

So, freezing point is -3.534°C

 

Please hit Yes button if I helped